OpenStudy (vianne):

Which literal equation is equivalent to y = 0.4x? A. x = 0.4y B. y = 0.4/x C. 0.4 = xy D. 0.4 = y/x

OpenStudy (vianne):

Now, I chose C.

OpenStudy (anonymous):

Try it yourself. :P

OpenStudy (anonymous):

I mean try it yourself first. :P

OpenStudy (vianne):

Because x times x = crossed out And x times y = xy

OpenStudy (vianne):

While 0.4 becomes by itself.

OpenStudy (vianne):

Am I correct? ;)

OpenStudy (anonymous):

Nope, sorry.

OpenStudy (vianne):

Nuuu

OpenStudy (anonymous):

wanna try one more time?

OpenStudy (vianne):

You divide it?

OpenStudy (anonymous):

Just tell me are you dividing by \(x\) or multiplying by \(x\). ?

OpenStudy (vianne):

Multiplying by x

OpenStudy (anonymous):

Okay I show you when you will do that, what you would get. :P

OpenStudy (vianne):

Okay.

OpenStudy (anonymous):

\[y \color{red}{\times x} = 0.4x \color{red}{\times x}\]

OpenStudy (vianne):

Yeye that's what I did :3

OpenStudy (anonymous):

\[x \times x = ?\]

OpenStudy (vianne):

Can't find that cross symbol >.<

OpenStudy (anonymous):

\[x \times x = x^2 \quad \qquad \quad x \times \frac{1}{x} = 1\]

OpenStudy (vianne):

oh..

OpenStudy (anonymous):

To cross that, one \(x\) should be in multiplication and other in divide.

OpenStudy (vianne):

Ohhh.......... ,-,

OpenStudy (anonymous):

So, best option is to divide by \(x\) and not multiply.

OpenStudy (vianne):

Which would be.. y/x?

OpenStudy (anonymous):

\[\frac{y}{\color{green}{x}} = \frac{0.4x}{\color{green}{x}}\]

OpenStudy (anonymous):

\[\frac{y}{\color{green}{x}} = \frac{0.4x}{\color{green}{x}} = \frac{y}{x} = \frac{0.4 \cancel{x}}{\cancel{x}} \implies \frac{y}{x} = 0.4\]

OpenStudy (vianne):

ohk

OpenStudy (anonymous):

So, here only one option is correct, okay?

OpenStudy (vianne):

Yes, which is 0.4 = y/x

OpenStudy (anonymous):

Do you know how we decide that we have to multiply here or divide here ??

OpenStudy (vianne):

yep

OpenStudy (anonymous):

If you know then how come you multiplied by \(x\) and not divide? :P

OpenStudy (vianne):

heh.......

OpenStudy (vianne):

That was after you explained :3

OpenStudy (anonymous):

Good.. :) Keep on practising these things, you will surely get one day. :) Now which one day, I don't know, it depends on you. :P

OpenStudy (vianne):

mk..

OpenStudy (anonymous):

\(\color{blue}{\dagger}\).

OpenStudy (vianne):

Lemme try answering this question..

OpenStudy (anonymous):

this? you have answered it already. :P

OpenStudy (vianne):

Solve for m. E = cm/r A. m = E + r/c B. m = Er/c C. m = E - rc m = E – rc D. m = E + rc

OpenStudy (vianne):

okais.. SOO.. ;-;..

OpenStudy (anonymous):

So ?

OpenStudy (vianne):

-.- I don't get this question....

OpenStudy (anonymous):

Ha ha ha.. :) This is the same like previous. :P

OpenStudy (vianne):

I know, but.... It's like containing + and - and and... more more.. ;-;

OpenStudy (vianne):

B since it's containing division/

OpenStudy (anonymous):

Why are you looking at your options?

OpenStudy (anonymous):

Suppose you are not given with options. yes by options, you can clearly find what will be your answer, but don't go with options, solve it step by step.

OpenStudy (anonymous):

B is surely your answer, as you said right.

OpenStudy (vianne):

E = cm/r So the first step is... Getting m = [something] right?

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

Cancel them one by one, let us cancel \(r\) first, okay?

OpenStudy (anonymous):

to cross out \(r\) from right side, what will you go?

OpenStudy (vianne):

okay

OpenStudy (anonymous):

*do?

OpenStudy (vianne):

e = cm

OpenStudy (vianne):

er*

OpenStudy (anonymous):

Yeah it is right, but e should be \(E\), capital one.

OpenStudy (vianne):

Why does it make a dif? o_o

OpenStudy (anonymous):

Now let us break the legs of c, because it is the option you chose for your previous question, and it made your answer incorrect. :P

OpenStudy (anonymous):

How can you remove \(c\) from right side?

OpenStudy (vianne):

um..

OpenStudy (anonymous):

Then do it.

OpenStudy (vianne):

c divided by cm?

OpenStudy (anonymous):

yep..

OpenStudy (vianne):

which would be m and Er/c?

OpenStudy (anonymous):

In \(cm\), \(c\) is in multiplication with \(m\) like this : \(c \times m\), as \(c\) is in multiplication, so you will divide it by \(c\), to cancel it.

OpenStudy (vianne):

ok

OpenStudy (anonymous):

\[m = \frac{Er}{c}\]

OpenStudy (anonymous):

How many left now?

OpenStudy (vianne):

For?

OpenStudy (anonymous):

I mean how many more problem you are left with?

OpenStudy (vianne):

Nonee.. I did 2 problems by myself and got 100% :3

OpenStudy (vianne):

and 3 problems with your help. (which was towards the end)

OpenStudy (anonymous):

So, anything more I can do it for you now?

OpenStudy (vianne):

nope.

OpenStudy (anonymous):

Satisfied?